# TurkicNLP: An NLP Toolkit for Turkic Languages

### v1

@@ 1.3 Contributions
  • Multi-Backend Integration: Rule-based Apertium FST morphological analyzers for 20 languages and Stanza neural models for the 5 languages with Universal Dependencies treebanks, selectable without API changes. NLLB-200 provides cross-lingual embeddings and translation for all 24 languages.
@@ 2.2 Turkic-Specific NLP Tools and Pipelines
Turkish: Zemberek-NLP Akın and Akın (2007) is a comprehensive Java library for Turkish NLP covering morphological analysis, POS tagging, and spell checking; it is restricted to Turkish, does not expose a pip-installable Python API, and has no cross-family support. TurkishDelightNLP Alecakir et al. (2022) is a neural Turkish NLP pipeline offering tokenization, morphological analysis, dependency parsing, and NER with strong performance on Turkish; it targets Turkish exclusively and does not integrate with other Turkic languages or script conversion. VNLP Turker et al. (2023) is an open-source Python NLP library for Turkish providing morphological analysis, named entity recognition, dependency parsing, and sentiment analysis, including quantitative benchmarks on standard Turkish test sets; like TurkishDelightNLP it is Turkish-only and provides no script handling. Istanbul Technical University (ITU) [^1] https://bbf.itu.edu.tr/en/research/research-labs/nlp-lab maintains a complementary Turkish NLP infrastructure including neural pipelines, the IMST and ITU Web treebanks, TRopBank PropBank, and supporting corpora Eryiğit et al. (2023). The Turkish NLP Suite Demir (2023) is an open-source non-profit collection that includes spaCy-compatible Turkish models, the SentiTurca sentiment corpus, TrGLUE benchmark, InstructTurca instruction-tuning data, and BellaTurca.
@@
Kyrgyz: KyrgyzNER Turdubaeva et al. (2024) provides 10,900 annotated sentences across 27 entity classes for Kyrgyz NER.
Multiple languages: Apertium Forcada et al. (2011) provides rule-based morphological analyzers and machine translation for 20+ Turkic languages via finite-state transducers, representing the most comprehensive freely available resource for Turkic morphological analysis. Its native interface is command-line-based, loading Apertium FST binaries via the hfst Python package without requiring system-level installation. Abdurakhmonova et al. 2022 present an NLP tool for linguistic analysis of Turkic languages focusing on morphological annotation, though it addresses a limited language set without a Python package interface. Universal Dependencies treebanks for Turkic languages were created by dedicated community efforts: the Turkish IMST treebank Sulubacak et al. (2016), the Tatar NMCTT treebank Taguchi et al. (2022), the Uyghur UDT Eli et al. (2016), the Yakut/Sakha treebank Merzhevich and Ferraz Gerardi (2022), and Old Turkish Derin and Harada (2021), with a parallel aligned treebank for Azerbaijani, Kyrgyz, Turkish, and Uzbek Eslami et al. (2025). These contributions remain language-specific and are not integrated into a unified pipeline toolkit.
@@ 2.3 Multilingual Models and Benchmarks
Azerbaijani: aLLMA Hajili et al. (2024) is a language model for Azerbaijani.
@@
Evaluation benchmarks: Evaluation benchmarks for Turkic NLP have advanced considerably. TUMLU Isbarov et al. (2025) covers 9 Turkic languages with 38K multiple-choice questions; Kardeş-NLU Senel et al. (2024) assesses cross-lingual NLU transfer across Turkic languages; and TR-MMLU Bayram et al. (2025) provides 6,200 multiple-choice questions across 62 subject categories specifically for benchmarking Turkish LLMs. Veitsman and Hartmann 2025 survey the state of NLP for Central Asian Turkic languages (Kazakh, Kyrgyz, Uzbek, Uyghur, Turkmen), documenting persistent resource gaps and the challenges of cross-lingual transfer from higher-resource Turkic languages. Collectively, these benchmarks underscore the need for unified processing infrastructure to support training and evaluation pipelines.
@@ 2.4 Resource Collections and Community Initiatives
Several community-driven collections catalogue the growing body of Turkic NLP resources. The Turkish NLP resource list Turkish NLP Community (2023) [^3] https://github.com/maidis/nlp-tr is a community-maintained index of Turkish NLP tools, corpora, treebanks, and pre-trained models. The Turkic Interlingua (TIL) project Syzdykov and et al. (2021)  [^4] https://github.com/turkic-interlingua/til-mt provides parallel corpora across 22 Turkic languages together with baseline systems for machine translation, automatic speech recognition, and text-to-speech, enabling systematic multilingual research across the family. Organisational and workshop infrastructure complements these datasets: SIGTURK, the ACL Special Interest Group on Turkic NLP SIGTURK Organizing Committee (2026) [^5] https://sigturk.github.io, organises dedicated workshop venues that bring together researchers working on Turkic languages and provide an important venue for disseminating resources and evaluation results.
@@ 4.4 Language and Processor Coverage
Table 3 summarises the script and processor availability for all 24 supported languages. Coverage reflects the state of publicly available resources: Turkish and Kazakh receive the full pipeline including NER; five languages have Stanza neural parsing via their UD treebanks; 20 languages have Apertium FST morphological analysis; and all 24 languages receive NLLB-200 sentence embeddings and machine translation.
@@
Azerbaijani Latn/Cyrl $\times$ $\times$ $\times$ $\times$
Iranian Azerbaijani Arab $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Turkmen Latn/Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$
Gagauz Latn $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
@@
Tatar Cyrl/Latn $\times$ $\times$ $\times$ $\times$
Bashkir Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$
Crimean Tatar Latn/Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$
Karakalpak Latn/Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Nogai Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Kumyk Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Karachay-Balkar Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
@@
Uzbek Latn/Cyrl $\times$ $\times$ $\times$ $\times$
@@
Sakha Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Altai Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Tuvan Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Khakas Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
@@
Chuvash Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
@@
Old Turkish Orkh $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
@@
Khalaj Arab $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Table 3: Script and processor availability for all 24 TurkicNLP languages. Script codes: Latn = Latin, Cyrl = Cyrillic, Arab = Perso-Arabic, Orkh = Old Turkic Runic; primary script listed first for dual-script languages.  = production/stable, $\sim$ = beta/prototype, $\times$ = not available, – = not applicable.
@@ 5 Processing Capabilities
The processing pipeline maps each stage to one or more interchangeable backends. For tokenization, a rule-based segmenter handles Latin- and Cyrillic-script input, with a dedicated Arabic-script tokenizer for right-to-left languages; neural Stanza Qi et al. (2020) tokenizers are used for the five UD-treebank languages (Turkish, Kazakh, Kyrgyz, Uyghur, Ottoman Turkish). Morphological analysis draws on Apertium FST binaries Forcada et al. (2011) for 20 languages. POS tagging, lemmatization, and dependency parsing are provided via Stanza neural models for those five languages. Named entity recognition targets Turkish and Kazakh. Sentence embeddings and machine translation use NLLB-200 NLLB Team et al. (2022) for all 24 languages. Table 3 summarises component availability across all 24 languages; runnable code samples for all supported languages are available in the companion repository. [^6] https://github.com/turkic-nlp/turkic-nlp-code-samples The following subsections describe each stage in detail.

### v4

@@ 1.3 Contributions
  • Multi-Backend Integration: Rule-based Apertium FST morphological analyzers for 20 languages and Stanza neural models for 10 languages—five with Universal Dependencies treebanks and five with custom-trained models—selectable without API changes. NLLB-200 provides cross-lingual embeddings and translation for all 24 languages.
@@ 2.2 Turkic-Specific NLP Tools and Pipelines
Turkish: Oflazer and Saraçlar Oflazer and Saraçlar (2018) provide the definitive reference volume on Turkish NLP, surveying 25 years of work on morphological processing, language modeling, deep parsing, machine translation, treebanks, and lexical resources. Zemberek-NLP Akın and Akın (2007) is a comprehensive Java library for Turkish NLP covering morphological analysis, POS tagging, and spell checking; it is restricted to Turkish, does not expose a pip-installable Python API, and has no cross-family support. TurkishDelightNLP Alecakir et al. (2022) is a neural Turkish NLP pipeline offering tokenization, morphological analysis, dependency parsing, and NER with strong performance on Turkish; it targets Turkish exclusively and does not integrate with other Turkic languages or script conversion. VNLP Turker et al. (2023) is an open-source Python NLP library for Turkish providing morphological analysis, named entity recognition, dependency parsing, and sentiment analysis, including quantitative benchmarks on standard Turkish test sets; like TurkishDelightNLP it is Turkish-only and provides no script handling. Istanbul Technical University (ITU) [^1] https://bbf.itu.edu.tr/en/research/research-labs/nlp-lab maintains a complementary Turkish NLP infrastructure including neural pipelines, the IMST and ITU Web treebanks, TRopBank PropBank, and supporting corpora Eryiğit et al. (2023). The Turkish NLP Suite Altinok (2023) is an open-source non-profit collection that includes spaCy-compatible Turkish models, the SentiTurca sentiment corpus, TrGLUE benchmark, InstructTurca instruction-tuning data, and BellaTurca.
@@
Kyrgyz: Alekseev and Turatali (2025) survey the state of Kyrgyz NLP, cataloguing available resources and proposing a development roadmap; they note that despite millions of speakers, Kyrgyz remains severely under-resourced. KyrgyzNER Turatali et al. (2025a) provides 10,900 annotated sentences across 27 entity classes for Kyrgyz NER. Turatali et al. (2025b) present a benchmark for evaluating Kyrgyz LLMs covering commonsense reasoning and multilingual understanding tasks. The Cramer Project The Cramer Project (2024) provides dedicated Kyrgyz speech models: AkylAI-STT-small (Whisper fine-tuned for Kyrgyz ASR) and AkylAI-TTS-mini (Matcha-TTS trained on 13 hours of Kyrgyz speech).
Multiple languages: Apertium Forcada et al. (2011) provides rule-based morphological analyzers and machine translation for 20+ Turkic languages via finite-state transducers, representing the most comprehensive freely available resource for Turkic morphological analysis. Its native interface is command-line-based, loading Apertium FST binaries via the hfst Python package without requiring system-level installation. Abdurakhmonova et al. 2022 present an NLP tool for linguistic analysis of Turkic languages focusing on morphological annotation, though it addresses a limited language set without a Python package interface. Universal Dependencies treebanks for Turkic languages were created by dedicated community efforts: the Turkish IMST treebank Sulubacak et al. (2016), the Tatar NMCTT treebank Taguchi et al. (2022), the Uyghur UDT Eli et al. (2016), the Yakut/Sakha treebank Merzhevich and Ferraz Gerardi (2022), and Old Turkish Derin and Harada (2021), with a parallel aligned treebank for Azerbaijani, Kyrgyz, Turkish, and Uzbek Akhundjanova et al. (2025). These contributions remain language-specific and are not integrated into a unified pipeline toolkit. Mirzakhalov et al. (2021) provide the first large-scale MT study across 22 Turkic languages, contributing a parallel corpus of approximately 1.4 million sentences, bilingual baselines for 26 language pairs, multi-domain test sets, and human evaluation scores. Tukeyev et al. (2026) extend this line of work by developing and cleaning parallel corpora for six Turkic languages (Kazakh, Azerbaijani, Kyrgyz, Turkish, Turkmen, and Uzbek) totalling over 3.8 million sentences, and fine-tuning NLLB-200 to achieve substantial gains in BLEU and chrF over the baseline for low-resource pairs.
@@ 2.3 Multilingual Models and Benchmarks
Azerbaijani: aLLMA Isbarov et al. (2024) is a language model for Azerbaijani.
@@
Evaluation benchmarks: Evaluation benchmarks for Turkic NLP have advanced considerably. TUMLU Isbarov et al. (2025) covers 9 Turkic languages (Azerbaijani, Crimean Tatar, Turkish, Uyghur, Uzbek, Karakalpak, Kyrgyz, Kazakh, and Tatar) with 38K multiple-choice questions; Kardeş-NLU Senel et al. (2024) benchmarks cross-lingual NLU transfer from Turkish into five related languages (Azerbaijani, Kazakh, Kyrgyz, Uzbek, and Uyghur); TR-MMLU Bayram et al. (2025) provides 6,200 multiple-choice questions across 62 subject categories specifically for benchmarking Turkish LLMs; TurkishMMLU Yüksel et al. (2024) offers over 10,000 curriculum-based multiple-choice questions across 9 subjects spanning sciences, mathematics, and culturally specific topics for Turkish; Cetvel Abrek Er et al. (2025) is a unified 23-task evaluation suite assessing language understanding, generation, and cultural competence of LLMs in Turkish; and TurkBench Toraman et al. (2026) provides a broad multi-task benchmark for evaluating Turkish LLMs across diverse NLP tasks. Turatali et al. (2025b) build a dedicated benchmark for Kyrgyz LLMs targeting commonsense reasoning and multilingual transfer. Veitsman and Hartmann 2025 survey the state of NLP for Central Asian Turkic languages (Kazakh, Kyrgyz, Uzbek, Uyghur, Turkmen), documenting persistent resource gaps and the challenges of cross-lingual transfer from higher-resource Turkic languages. Collectively, these benchmarks underscore the need for unified processing infrastructure to support training and evaluation pipelines.
2.4 Speech Technologies for Turkic Languages
Speech processing resources for Turkic languages have expanded significantly through dedicated corpus construction and multilingual modeling efforts. ISSAI has led the development of both ASR and TTS systems spanning the Turkic family: TurkicASR Mussakhojayeva et al. (2023a) trains multilingual ASR models across 10 Turkic languages (Azerbaijani, Bashkir, Chuvash, Kazakh, Kyrgyz, Sakha, Tatar, Turkish, Uyghur, Uzbek), achieving a 56.7% relative CER reduction through cross-lingual transfer compared to monolingual baselines. TurkicTTS Yeshpanov et al. (2023) extends this multilingual paradigm to speech synthesis for a partially overlapping set of 10 languages (Azerbaijani, Bashkir, Kazakh, Kyrgyz, Sakha, Tatar, Turkish, Turkmen, Uyghur, Uzbek)—notably including Turkmen but not Chuvash. Söyle Mussakhojayeva et al. (2023b) provides noise-robust ASR for 11 Turkic languages by fine-tuning Whisper with data augmentation.
Key speech corpora include KSC2 (1,200 hours of Kazakh; Mussakhojayeva et al. 2021a), TSC (218 hours of Turkish; ISSAI 2024), USC (105 hours of Uzbek; Mussakhojayeva et al. 2021b), TatSC (269 hours of Tatar; ISSAI 2023), TurkmenSpeech (251 hours of Turkmen; Mamedov 2024), and KazakhTTS2 (271 hours; Mussakhojayeva et al. 2022). Mozilla Common Voice Ardila et al. (2020) provides community-recorded speech data for several Turkic languages, with Bashkir ($\sim$265 hours), Turkish ($\sim$129 hours), Kyrgyz ($\sim$44 hours), Tatar ($\sim$29 hours), Chuvash ($\sim$12 hours), Sakha ($\sim$6 hours), and Kazakh ($\sim$1 hour). FLEURS Conneau et al. (2022) covers Azerbaijani, Kazakh, Kyrgyz, Turkish, and Uzbek for few-shot speech evaluation. Dedicated Kyrgyz speech models have also emerged: the Cramer Project The Cramer Project (2024) releases AkylAI-STT-small (Whisper fine-tuned for Kyrgyz) and AkylAI-TTS-mini (Matcha-TTS trained on 13 hours of Kyrgyz speech), providing the first openly available language-specific ASR and TTS systems for Kyrgyz.
2.5 Tokenizer Efficiency Across Turkic Languages
(a) Mean fertility (tokens per word).
(b) 95th-percentile tokens per word.
Figure 2: Tokenizer efficiency across 20 Turkic languages and English for four LLM tokenizers: Qwen3.5, LLaMA 3, OLMo 3, and GPT-5.2. Higher fertility indicates less efficient tokenization, reflecting poorer vocabulary coverage for agglutinative languages.
The choice of tokenizer has a profound impact on inference cost and downstream performance for Turkic languages. We evaluate four widely used LLM tokenizers—Qwen3.5-35B, Meta-LLaMA-3-8B, OLMo-3-1025-7B, and GPT-5.2—across 20 Turkic languages using a parallel set of 20 typologically diverse sentences translated into each language (Figure 2).
English exhibits near-optimal fertility of 1.08–2.09 tokens per word across all tokenizers. Turkish, the best-represented Turkic language, reaches 1.95 (Qwen/GPT-5.2) to 3.03 (LLaMA). For under-represented languages, the “tokenizer tax” is severe: Uyghur fertility ranges from 3.38 (GPT-5.2) to 6.85 (OLMo), Sakha from 3.48 (GPT-5.2) to 5.21 (LLaMA), and Bashkir from 2.54 (GPT-5.2) to 5.68 (LLaMA). The 95th-percentile tokens per word—a proxy for worst-case segmentation of complex agglutinated forms—reaches 11.0 for Uyghur and Tatar under LLaMA, compared to 2.0 for English across all tokenizers.
GPT-5.2 and Qwen3.5 consistently achieve the best Turkic coverage, while LLaMA 3 exhibits the highest fertility across all Turkic languages. These findings underscore the need for tokenizer-aware evaluation when deploying multilingual LLMs on Turkic text and motivate continued development of dedicated Turkic tokenization strategies.
@@ 2.6 Resource Collections and Community Initiatives
Several community-driven collections catalogue the growing body of Turkic NLP resources. The Turkish NLP resource list Turkish NLP Community (2023) [^3] https://github.com/maidis/nlp-tr is a community-maintained index of Turkish NLP tools, corpora, treebanks, and pre-trained models. The Turkic Interlingua (TIL) project Syzdykov and et al. (2021)  [^4] https://github.com/turkic-interlingua/til-mt provides parallel corpora across 22 Turkic languages together with baseline systems for machine translation, automatic speech recognition, and text-to-speech, enabling systematic multilingual research across the family. The UD-Turkic research group [^5] https://github.com/ud-turkic works on harmonising Universal Dependencies treebanks across Turkic languages, producing the first fully aligned parallel UD treebanks for Azerbaijani, Kyrgyz, Turkish, and Uzbek Akhundjanova et al. (2025). Organisational and workshop infrastructure complements these datasets: SIGTURK, the ACL Special Interest Group on Turkic NLP SIGTURK Organizing Committee (2026) [^6] https://sigturk.github.io, organises dedicated workshop venues that bring together researchers working on Turkic languages and provide an important venue for disseminating resources and evaluation results. The TurkLang conference series [^7] https://turklang.ieees.org/, held annually since 2013, serves as another key venue dedicated to computational processing of Turkic languages, fostering cross-institutional collaboration across the Turkic-speaking world. Cekinel and Çöltekin (2025) provide a recent systematic survey of Turkic language modeling, reviewing available corpora, models, and benchmarks while highlighting persistent challenges around data scarcity, tokenization, and multilingual transfer across the family.
@@ 4.4 Language and Processor Coverage
Table 3 summarises the script and processor availability for all 24 supported languages. Coverage reflects the state of publicly available resources: Turkish and Kazakh receive the full pipeline including NER; ten languages have Stanza neural parsing via UD treebanks (five official, five custom-trained); 20 languages have Apertium FST morphological analysis; a multilingual Glot500-based neural morphological analyzer and lemmatizer covers 20 languages; and all 24 languages receive NLLB-200 sentence embeddings and machine translation.
@@
Azerbaijani Latn/Cyrl $\times$
South Azerb. Arab $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Turkmen Latn/Cyrl $\sim$ $\times$
Gagauz Latn $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$ $\times$
@@
Tatar Cyrl/Latn $\times$
Bashkir Cyrl $\sim$ $\times$ $\times$
Crimean Tatar Latn/Cyrl $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$
Karakalpak Latn/Cyrl $\sim$ $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$
Nogai Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Kumyk Cyrl $\sim$ $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$ $\times$
Karachay-Balkar Cyrl $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
@@
Uzbek Latn/Cyrl $\times$
@@
Sakha Cyrl $\sim$ $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$ $\times$
Altai Cyrl $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$
Tuvan Cyrl $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$
Khakas Cyrl $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$
@@
Chuvash Cyrl $\sim$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$
@@
Old Turkish Orkh $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
@@
Khalaj Arab $\times$ $\sim$ $\sim$ $\sim$ $\times$ $\times$ $\times$ $\times$ $\times$
Table 3: Script and processor availability for all 24 TurkicNLP languages. Script codes: Latn = Latin, Cyrl = Cyrillic, Arab = Perso-Arabic, Orkh = Old Turkic Runic; primary script listed first for dual-script languages. Morph (Neu) = Glot500-based neural morphological analyzer and lemmatizer. POS Tag, Lemmatize, and Dep. Parse reflect combined Stanza and Glot500 backend coverage.  = production/stable, $\sim$ = beta/prototype/zero-shot, $\times$ = not available, – = not applicable.
@@ 5 Processing Capabilities
The processing pipeline maps each stage to one or more interchangeable backends. For tokenization, a rule-based segmenter handles Latin- and Cyrillic-script input, with a dedicated Arabic-script tokenizer for right-to-left languages; neural Stanza Qi et al. (2020) tokenizers are used for the five UD-treebank languages (Turkish, Kazakh, Kyrgyz, Uyghur, Ottoman Turkish) and for five additional languages with custom-trained models (Uzbek, Azerbaijani, Turkmen, Tatar, Bashkir). Morphological analysis draws on Apertium FST binaries Forcada et al. (2011) for 20 languages. POS tagging, lemmatization, and dependency parsing are provided via Stanza neural models for those ten languages. Two multilingual neural models built on a frozen Glot500 Imani et al. (2023) backbone extend neural POS tagging and dependency parsing to 15 languages and morphological analysis with lemmatization to 23 languages. Named entity recognition targets Turkish and Kazakh. Sentence embeddings and machine translation use NLLB-200 NLLB Team et al. (2022) for all 24 languages. Table 3 summarises component availability across all 24 languages; runnable code samples for all supported languages are available in the companion repository. [^8] https://github.com/turkic-nlp/turkic-nlp-code-samples The following subsections describe each stage in detail.